OpenStudy (anonymous):

differentiate cot^2(3x+4) I'm assuming chain rule? But I can't get the steps to work out and get a nice, simplified answer

hartnn (hartnn):

chain rule is correct, you should know derivative of x^2 , cot x and 3x+4 to solve this. if you have tried, show us what u get...

OpenStudy (anonymous):

Okay, so 2x(-csc^2)(3x+4)(3)?

hartnn (hartnn):

using chain rule, derivative of \((f(x))^2\) will be 2f(x)f'(x) right ? you are missing the f(x) part there.

OpenStudy (anonymous):

oh, no. Okay, let me try again. 2(cotx(3x+4))(3)

OpenStudy (anonymous):

Wait wait wait. . . ugh. I know I'm getting the middle part incorrectly

hartnn (hartnn):

take you time, i can wait. try to get to correct answer, i will correct you incase of any error.

OpenStudy (anonymous):

so f(x) = cot^2(3x+4), which means that is equivalent to cot(3x+4)^2. then 2(cot(3x+4))(-3csc^2)?

hartnn (hartnn):

getting very close to correct answer! you basically missed the argument of csc^2 shouldn't it be, 2(cot(3x+4))(-3csc^2(3x+4)) ? and with one simple step of simplification, you'll get the required answer.

OpenStudy (anonymous):

Oh, I see. I just derived the 3x+4 from the second part into "-3" and totally left out the rest. Then,is the final answers -6cosssin(3x+4)? How do I put the cos(3x+4) and the sin(3x+4) together to pull out the "(3x+4)"? @hartnn

hartnn (hartnn):

what?! i couldn't understand your question, if you got this : 2(cot(3x+4))(-3csc^2(3x+4)) then its simplified version will be -6(cot(3x+4))csc^2(3x+4) i believe this is simplified and needs no further simplification

OpenStudy (anonymous):

ohh, okay. No, that is what I got. But I thought since it had cot(3x+4) and csc^2(3x+4), maybe there was some way I can take out the (3x+4) since it was in both But if that's as simplified as it can get, then I see I can't just pull it out

hartnn (hartnn):

ok, that (3x+4) is the angle (or argument) of the trigonometric function. like \(\theta\) so, say in \(\cos \theta \sin^2 \theta\), you cannot factor \(\theta\) and just write cos . sin^2 , in same way you cannot factor out 3x+4 here.

OpenStudy (anonymous):

OH, wow. I see, I see. that makes sense, thank you.

hartnn (hartnn):

welcome ^_^